# A Disentangled Representation Learning Framework for Low-altitude Network Coverage Prediction

### v1

@@ VI Conclusion
Our research has introduced two pioneering techniques: feature compression based on expert knowledge and a model-guided disentangled representation network. The former not only achieves data compression but also addresses the issue of feature imbalance, while the latter, by integrating a signal propagation model, mitigates reliance on extensive data and reduces the risk of overfitting. Both of the techniques have demonstrated their efficacy in our experimental validation based on real-world data. These methodologies significantly enhance the accuracy of low-altitude coverage prediction.
However, we acknowledge the limitations of our study, particularly the insufficient consideration of environmental obstructions affecting signal propagation in low-altitude networks, and the opportunity to integrate more precise propagation models into our neural network architecture. Future endeavors will focus on exploiting large-scale data and large language models to refine LANC predictions. We also aim to incorporate detailed environmental data and BS beamforming patterns into our models, thus improving the precision and resilience of LANC.

### v2

@@ VI Conclusion
Our research has introduced two pioneering techniques: feature compression based on expert knowledge and a model-guided disentangled representation network. The former not only achieves data compression but also addresses the issue of feature imbalance, while the latter, by integrating a signal propagation model, mitigates reliance on extensive data and reduces the risk of overfitting. Both of the techniques have demonstrated their efficacy in our experimental validation based on real-world data, significantly enhancing the accuracy of low-altitude coverage prediction.
However, we acknowledge the limitations of our study, particularly the insufficient consideration of environmental obstructions affecting signal propagation in low-altitude networks. According to related works, incorporating environmental information could potentially improve prediction accuracy by approximately 0.5-1 $\mathrm{dB}$ RMSE for radio map reconstruction problems [16], and by 1-2 $\mathrm{dB}$ RMSE for radio map prediction without sparse sampling [40], where the larger gain is attributed to the absence of measurement data. These quantitative insights highlight the potential impact of environmental factors on prediction performance. Additionally, there remains an opportunity to integrate more precise propagation models into our neural network architecture.
Future endeavours will focus on exploiting large-scale data and large language models to refine LANC predictions. We also aim to incorporate detailed environmental data and BS beamforming patterns into our models, thus improving the precision and resilience of LANC while addressing the identified limitations.